OpenStudy (anonymous):

whats the area of the circle d=12ft

OpenStudy (nubeer):

radius = d/2 area of circle. = pi*r^2

OpenStudy (nathan917):

The area (in square footage) of a circle with a diameter of 12 feet can be calculated a few ways: 1. Area = pi (3.1416) times r^2 2. Area = pi (3.1416) times (d/2)^2 where d = the diameter = 2 radii 3 Area = pi (3.1416) times d^2/4 where d = diameter The reason that number 3 method is so neat is that pi (3.1416) divided by 4 becomes a constant (0.7854) and when you multiply it by the diameter squared, you are performing the same operation as number 1 and number 2 method much quicker. If you can remember the constant (0.7854) or the method used to calculate the constant (pi (3.1416) divided by 4 = 0.7854), this works for any diameter circle. For example, determine the area of a circle with a 12 foot diameter: Area = 0.7854 times d^2 = 0.7854(12 feet)^2 = 0.7854 times 144 sq. ft. = 113.0976 square feet

OpenStudy (nubeer):

lol copy paste. :P

OpenStudy (nathan917):

Yes. Anything wrong with that?

OpenStudy (anonymous):

36ft?

OpenStudy (nubeer):

yeah just too long to read it.

OpenStudy (nathan917):

113.09 Ft.

OpenStudy (anonymous):

thats wrong:(

OpenStudy (nathan917):

For example, determine the area of a circle with a 12 foot diameter: Area = 0.7854 times d^2 = 0.7854(12 feet)^2 = 0.7854 times 144 sq. ft. = 113.0976 square feet

OpenStudy (nathan917):

Questions?

OpenStudy (nubeer):

how about 113.04 sq. ft

OpenStudy (nathan917):

Yes....

OpenStudy (nathan917):

Love said Ft so i said Ft i was going to say that though..